F. Frans is the workshop supervisor of NBW. She is paid a fixed salary monthly of N$10000. To keep her motivated she is entitled to an annual bonus of N$9000 to her pension fund annually. She works 9 hours per day from Monday to Friday every week. She is also entitled to 28 days paid vacation leave. There are 12 public holidays in the year for which 9 fall on weekends. Assume a 365 days year.

Required:

Using information relating to F. Frans compute Frans s hourly recovery tariff per hour.


1
Expert's answer
2022-03-17T09:39:17-0400

The year has 365 days

"\frac{365}{7} =" 52 weeks and 1 day


Works 5 days every week

Therefore "(52\u00d7 5)+1 = 261" days

The 28 vacation leave days are accounted above.


If Frans works 9 hours a day

The the total hours annually will be

"261\u00d79 =2349"

However the public holidays will be deducted since there is no information on how Frans is remunerated.

Thus "(261-3)\u00d79 = 2322" hours

The hourly recovery tariff is

"\frac{Total. pay}{total. hours}"


Total pay = "10000\u00d712 = 120000 + 9000" = 129000

Therefore hourly recovery tariff is;

"\frac{129000}{2322}= N 55.556"